a water solution contains 10% ethyl alcohol by mass and has a density of 0.983 g/cm3 . what is the mass of ethyl alcohol is present in 7.5 L of this solution
1
Expert's answer
2016-10-13T14:45:03-0400
7.5L = 7.5*1000 mL
mass of 7.5 L solution: 7.5*1000 mL * 0.983 g/cm 3 = 7.3725*1000 g
mass of ethyl alcohol: 7.3725*1000 g * 0.1 = 737.25 g

Answer: mass of ethyl alcohol is 737.25 g.